OpenStudy (anonymous):

Multiply these polynomials. (4x2 - 3x)(5x2 - 6x + 2) A. 20x4 - 39x3 + 26x2 - 6x B. 20x4 - 39x3 + 18x2 - 6x C. 20x4 - 24x3 + 18x2 - 6x D. 20x4 - 24x3 + 26x2 - 6x

OpenStudy (anonymous):

20x4 - 24x3 + 18x2 - 6x

OpenStudy (hoblos):

(4x2 - 3x)(5x2 - 6x + 2) = 4x2(5x2 - 6x + 2) - 3x(5x2 - 6x + 2) = 20x4 -24x3 +8x2 -15x3 +18x2 -6x = 20x4 - 39x3 + 26x2 - 6x